What is the correct way of writting 1970 in Roman Numeral pls tell
Found 2 solutions by Alan3354, CoolRekt:
Answer by Alan3354(69443)

What is the correct way of writting 1970 in Roman Numeral pls tell
MCMLXX

Answer by CoolRekt(53)

MCMLXX = 1970
M = 1000
CM = 900
L = 50
XX = 20
M + CM + L + XX = 1000 + 900 + 50 + 20 = MCMLXX = 1970
CoolRekt